OpenStudy (anonymous):

What are the coordinates of the vertex of the graph of y= x^2+6x-11? A) (3, -2) B) (3, 16) C) (-3, -29) D) (-3, -20)

OpenStudy (anonymous):

The x-coordinate 0f the vertex( or the axis of the parabola) x = -b/2a = -6/2 = -3 The y-coordinate of the vertex: y(3) = (-3)^2 + 6(-3) - 11 = .......

OpenStudy (anonymous):

@thu1935 -20?

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

y(3) = (-3)^2 + 6(-3) - 11 = 9 - 18 - 11

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

nvm, -20 is right

ganeshie8 (ganeshie8):

good job :)

OpenStudy (anonymous):

ok. thanks @ganeshie8!

ganeshie8 (ganeshie8):

u wlc !!